# Swin-UMamba: Mamba-based UNet with ImageNet-based pretraining

### v1

@@ Abstract
Accurate medical image segmentation demands the integration of multi-scale information, spanning from local features to global dependencies. However, it is challenging for existing methods to model long-range global information, where convolutional neural networks (CNNs) are constrained by their local receptive fields, and vision transformers (ViTs) suffer from high quadratic complexity of their attention mechanism. Recently, Mamba-based models have gained great attention for their impressive ability in long sequence modeling. Several studies have demonstrated that these models can outperform popular vision models in various tasks, offering higher accuracy, lower memory consumption, and less computational burden. However, existing Mamba-based models are mostly trained from scratch and do not explore the power of pretraining, which has been proven to be quite effective for data-efficient medical image analysis. This paper introduces a novel Mamba-based model, Swin-UMamba, designed specifically for medical image segmentation tasks, leveraging the advantages of ImageNet-based pretraining. Our experimental results reveal the vital role of ImageNet-based training in enhancing the performance of Mamba-based models. Swin-UMamba demonstrates superior performance with a large margin compared to CNNs, ViTs, and latest Mamba-based models. Notably, on AbdomenMRI, Encoscopy, and Microscopy datasets, Swin-UMamba outperforms its closest counterpart U-Mamba by an average score of 3.58%. The code and models of Swin-UMamba are publicly available at: https://github.com/JiarunLiu/Swin-UMamba.
1 Introduction
@@
There are several challenges that need to be addressed. First, given that the application of Mamba blocks in the vision domain is relatively new, further experimental validation is required to evaluate their performance and identify the most suitable structure of Mamba blocks for medical image segmentation tasks. Another challenge lies in the fact that existing Mamba-based models for medical image segmentation have not taken into account the transferability from ImageNet pretrained models. Consequently, the network structure requires redesigning to integrate the pretrained model. Third, there is a need for the scalability and efficiency of Mamba-based models for real-world deployment [31], particularly in resource-constrained environments, which is commonly found in medical practice.
@@
To the best of our knowledge, we are the first attempt to discover the impact of pretrained Mamba-based networks in medical image segmentation task. Our experiment verified that ImageNet-based pretraining plays an important role in medical image segmentation for Mamba-based networks, which sometimes is crucial.
@@ 2.3 Swin-UMamba decoder
We follow the commonly used U-shaped architecture with dense skip connections to construct Swin-UMamba. U-Net and its variations have demonstrated remarkable efficiency in medical image segmentation tasks. This architecture leverages skip connections for the recovery of low-level details and employs an encoder-decoder structure for high-level information extraction. To enhance the native up-sampling block in U-Net, we introduce two modifications: 1) an extra convolution block with a residual connection to process skip connection features, and 2) an additional segmentation head at each scale for deep supervision [33].
The structure of the up-sampling block was illustrated in Fig. 1. Given skip-connected feature from stage- and feature from the last up-sampling block, the output feature of -th up-sampling block and the segmentation map at stage- can be formulated as follows:
where , , are the feature concatenation operation, transpose convolution, and a segmentation head with convolution that project feature from dimension to the number of class , respectively. and are the height and width of the feature map at stage-. and are two convolution blocks with residual connection at stage-, each was composed of two convolution layers with LeakyRELU activation. In addition to the skip connections between encoding stages and up-sampling blocks, we add an extra skip connection from the input with operations. We use a convolution to get the final segmentation output.
@@ 2.4 Swin-UMamba: Swin-UMamba with Mamba-based decoder
Several modifications were made on Swin-UMamba. First, the up-sample blocks in Swin-UMamba were replaced by patch expanding [34] and two VSS blocks. We found that many parameters and computation burdens were caused by the heavy CNN-based decoder. Second, we use a patch embedding layer that directly projects input image from into feature maps of shape follow VMamba [24]. Notably, the last patch expanding block in Swin-UMamba is a up-sample operation, mirroring the patch embedding layer. The residual patch expanding layers were up-sampling operation. Skip connections originating from the input image and down-sampled features in Swin-UMamba were removed, as there were no corresponding decoding blocks for them. Additionally, deep supervision was applied at resolutions of , , , and , incorporating additional segmentation heads (i.e., convolutions mapping high-dimensional features to ) for each scale. Combining all these modifications, the number of network parameters was reduced from 40M to 27M, and the FLOPs were decreased from 58.4G to 15.0G on the AbdomenMRI dataset. Further statistics regarding the number of network parameters and FLOPs are provided in Table 1, Table 2, and Table 3. The structure of Swin-UMamba is illustrated in Fig. 2.
@@ 3.2 Implemetation details
We implemented Swin-UMamba on top of the well-established nnU-Net framework [11]. It’s self-configuring feature enabled us to focus on network design rather than other trivial details. The loss function is the sum of Dice loss and cross-entropy loss and we perform deep supervision [33] at each scale. We use an AdamW optimizer with weight decay following [24]. A cosine learning rate decay was adopted with an initial learning rate equal to 0.0001. We use the pretrained VMamba-Tiny model111The pretrained model can be found at: https://github.com/MzeroMiko/VMamba to initialize Swin-UMamba for all datasets. During training, we froze parameters from the pretrained model for the first 10 epochs to align other modules. Hyperparameters were kept consistent across all three datasets, except for the number of training epochs and data-specific settings (e.g., image patch size). Swin-UMamba was trained for 100 epochs on the AbdomenMRI dataset, 250 epochs on the Endoscopy dataset, and 350 epochs on the Microscopy dataset, respectively. Following [23], we disabled the testing time argumentation for a more streamlined and efficient evaluation. It’s worth noting that further improvement is possible with additional training and proper hyperparameter tuning, which we leave for future work. Our primary goal is to assess the impact of the pretrained models on medical image segmentation rather than solely aiming for state-of-the-art (SOTA) performance. For more details, please refer to our code implementation222https://github.com/JiarunLiu/Swin-UMamba.
@@ 3.3 Baselines and evaluation metrics
| Swin-UMamba | 27M | 15.0G | 0.6653 | 0.7312 |
| Swin-UMamba | 40M | 58.4G | 0.7464 | 0.8023 |
@@
| Swin-UMamba | 27M | 15.0G | 0.7705 | 0.8376 |
| Swin-UMamba | 40M | 58.4G | 0.7768 | 0.8442 |
@@ 3.4 Comparisons on AbdomenMRI dataset
Table 1 presents the segmentation performance on the AbdomenMRI dataset. Both Swin-UMamba and Swin-UMamba outperform all baseline methods, including CNN-based networks, transformer-based networks, and Mamba-based networks. The superior result demonstrates the great potential of the Mamba-based network in medical image segmentation. Notably, all Mamba-based networks outperform CNN-based and transformer-based baselines by at least 1% on both DSC and NSD. Swin-UMamba exhibits a remarkable 1.43% improvement in DSC over U-Mamba_Enc, which is the previous SOTA model on this dataset. As illustrated in Fig. 3a, Swin-UMamba can accurately recognize the shape and type of target organs, whereas U-Mamba_Enc failed to identify the type of all organs.
ImageNet-based pretraining plays a crucial role in our experiments, leading to a significant 3.04% improvement in DSC and a notable 4.19% improvement in NSD for Swin-UMamba. Moreover, leveraging ImageNet-based pretraining facilitates faster and more stable training, requiring merely one-tenth of the training iterations compared to baseline methods. A drastic phenomenon is observed with Swin-UMamba. Without ImageNet-based pretraining, Swin-UMamba fails to converge properly on this dataset with default settings. To address this issue, we disable the deep supervision of Swin-UMamba and extend its training epochs to 200. Despite that, Swin-UMamba outperforms all baseline methods when utilizing the ImageNet pretrained weights. This improvement is particularly noteworthy considering that Swin-UMamba has less than half of the network parameters and approximately one-third of the FLOPs compared to the previous SOTA model U-Mamba.
We also observed a disparity in parameter numbers and FLOPs between Swin-UMamba and Swin-UMamba. This discrepancy is primarily attributed to the CNN-based decoder, as Swin-UMamba and Swin-UMamba share almost identical structures in the encoder part. We opted to retain the CNN-based decoder because it did take better results in most cases.
@@
| Swin-UMamba | 28M | 35.9G | 0.6175 | 0.6320 |
| Swin-UMamba | 40M | 140.6G | 0.5466 | 0.5605 |
@@
| Swin-UMamba | 28M | 35.9G | 0.6717 | 0.6871 |
| Swin-UMamba | 40M | 140.6G | 0.6774 | 0.6926 |
@@ 3.5 Comparisons on Endoscopy dataset
Table 2 presents the segmentation performance of each model on the Endoscopy dataset. Swin-UMamba outperforms U-Mamba_Bot over 2.24% on DSC and 2.35% on NSD. Notably, the performance gap of using the pretrained model is substantially higher compared to the AbdomenMRI dataset. The visualized result of Swin-UMamba on Endoscopy is shown in Fig. 3b. Moreover, we observed an impressive performance gain of 13.08% on DSC and 13.21% on NSD with the pretrained model for Swin-UMamba. One possible explanation is that the Endoscopy dataset is smaller than the AbdomenMRI dataset, and models are prone to overfitting to the training data. Leveraging the power of a pretrained model is an effective strategy for mitigating overfitting in such small datasets. In addition, we found that Swin-UMamba performs better than Swin-UMamba when trained from scratch, possibly benefiting from its fewer parameters to avoid overfitting.
@@
| Swin-UMamba | 27M | 38.2G | 0.5153 |
| Swin-UMamba | 40M | 149.9G | 0.4833 |
@@
| Swin-UMamba | 27M | 38.2G | 0.5850 |
| Swin-UMamba | 40M | 149.9G | 0.5836 |
@@ 3.6 Comparisons on Microscopy dataset
Table 3 presents the segmentation performance on the Microscopy dataset. Swin-UMamba and Swin-UMamba continue to outperform all baseline methods by margins ranging from 2.42% to 18.83%. In contrast to previously mentioned datasets, the Microscopy dataset features higher image resolution, fewer samples, and greater visual difference. This imposes greater demands on the model’s capacity for long-range information modeling and data-efficiency. As shown in Fig. 3c, Swin-UMamba can accurately segment target cells while U-Mamba_Enc missing some. Somewhat surprisingly, we observe that Swin-UMamba outperforms Swin-UMamba on this dataset, indicating the potential of the pure Mamba-based network in settings with higher image resolution and limited samples.

### v2

@@ Abstract
Accurate medical image segmentation demands the integration of multi-scale information, spanning from local features to global dependencies. However, it is challenging for existing methods to model long-range global information, where convolutional neural networks (CNNs) are constrained by their local receptive fields, and vision transformers (ViTs) suffer from high quadratic complexity of their attention mechanism. Recently, Mamba-based models have gained great attention for their impressive ability in long sequence modeling. Several studies have demonstrated that these models can outperform popular vision models in various tasks, offering higher accuracy, lower memory consumption, and less computational burden. However, existing Mamba-based models are mostly trained from scratch and do not explore the power of pretraining, which has been proven to be quite effective for data-efficient medical image analysis. This paper introduces a novel Mamba-based model, Swin-UMamba, designed specifically for medical image segmentation tasks, leveraging the advantages of ImageNet-based pretraining. Our experimental results reveal the vital role of ImageNet-based training in enhancing the performance of Mamba-based models. Swin-UMamba demonstrates superior performance with a large margin compared to CNNs, ViTs, and latest Mamba-based models. Notably, on AbdomenMRI, Encoscopy, and Microscopy datasets, Swin-UMamba outperforms its closest counterpart U-Mamba_Enc by an average score of 2.72%. The code and models of Swin-UMamba are publicly available at: https://github.com/JiarunLiu/Swin-UMamba.
Keywords:
Medical image segmentation ImageNet-based pretraining Long-range dependency modeling.1 Introduction
@@
There are several challenges that need to be addressed. First, existing Mamba-based models for medical image segmentation have not taken into account the transferability from ImageNet pretrained models. Consequently, the network structure requires redesigning to integrate the pretrained model. Given the fact that the application of Mamba blocks in the vision domain is relatively new, further experimental evaluation is required for medical image segmentation tasks. Third, there is a need for the scalability and efficiency of Mamba-based models for real-world deployment [31], particularly in resource-constrained environments, which is commonly found in medical practice.
@@
To the best of our knowledge, we are the first attempt to discover the impact of pretrained Mamba-based networks in medical image segmentation. Our experiment verified that ImageNet-based pretraining plays an important role in medical image segmentation for Mamba-based networks, which sometimes is crucial.
@@ 2.3 Swin-UMamba decoder
We follow the commonly used U-shaped architecture with dense skip connections to construct Swin-UMamba. U-Net and its variations have demonstrated remarkable efficiency in medical image segmentation tasks. This architecture leverages skip connections for the recovery of low-level details and employs an encoder-decoder structure for high-level information extraction. To enhance the native up-sample block in U-Net, we introduce two modifications: 1) an extra convolution block with a residual connection to process skip connection features, and 2) an additional segmentation head at each scale for deep supervision [33].
The structure of the up-sample block was illustrated in Fig. 1. Given skip-connected feature from stage- and feature from the last up-sample block, the output feature of -th up-sample block and the segmentation map at stage- can be formulated as follows:
where , , are the feature concatenation operation, transpose convolution, and a segmentation head with convolution that project feature from dimension to the number of class , respectively. and are the height and width of the feature map at stage-. and are two convolution blocks with residual connection at stage-, each was composed of two convolution layers with LeakyRELU activation. In addition to the skip connections between encoding stages and up-sample blocks, we add an extra skip connection from the input with operations. We use a convolution to get the final segmentation output.
@@ 2.4 Swin-UMamba: Swin-UMamba with Mamba-based decoder
Several modifications were made on Swin-UMamba. First, the up-sample blocks in Swin-UMamba were replaced by patch expanding [34] and two VSS blocks. We found that many parameters and computation burdens were caused by the heavy CNN-based decoder. Second, we use a patch embedding layer that directly projects input image from into feature maps of shape follow VMamba [24]. Notably, the last patch expanding block in Swin-UMamba is a up-sample operation, mirroring the patch embedding layer. The residual patch expanding layers were up-sampling operation. Skip connections originating from the input image and down-sampled features in Swin-UMamba were removed, as there were no corresponding decoding blocks for them. Additionally, deep supervision was applied at resolutions of , , , and , incorporating additional segmentation heads (i.e., convolutions mapping high-dimensional features to ) for each scale. Combining all these modifications, the number of network parameters was reduced from 60M to 28M, and the FLOPs were decreased from 68.0G to 18.9G on the AbdomenMRI dataset. Further statistics regarding the number of network parameters and FLOPs are provided in Table 1, Table 2, and Table 3. The structure of Swin-UMamba is illustrated in Fig. 2.
@@ 3.2 Implemetation details
We implemented Swin-UMamba on top of the well-established nnU-Net framework [11]. It’s self-configuring feature enabled us to focus on network design rather than other trivial details. The loss function is the sum of Dice loss and cross-entropy loss and we perform deep supervision [33] at each scale. We use an AdamW optimizer with weight decay following [24]. A cosine learning rate decay was adopted with an initial learning rate . We use the pretrained VMamba-Tiny model111The pretrained model can be found at: https://github.com/MzeroMiko/VMamba to initialize Swin-UMamba for all datasets. During training, we froze parameters from the pretrained model for the first 10 epochs to align other modules. Hyperparameters were kept consistent across all three datasets, except for the number of training epochs and data-specific settings (e.g., image patch size). Swin-UMamba was trained for 100 epochs on the AbdomenMRI dataset, 350 epochs on the Endoscopy dataset, and 450 epochs on the Microscopy dataset, respectively. Following [23], we disabled the testing time argumentation for a more streamlined and efficient evaluation. It’s worth noting that further improvement is possible with additional training and proper hyperparameter tuning, which we leave for future work. Our primary goal is to assess the impact of the pretrained models on medical image segmentation rather than solely aiming for state-of-the-art (SOTA) performance. For more details, please refer to our code implementation222https://github.com/JiarunLiu/Swin-UMamba.
@@ 3.3 Baselines and evaluation metrics
Swin-UMamba | 60M | 68.0G | 0.7054 | 0.7647 |
Swin-UMamba | 28M | 18.9G | 0.6653 | 0.7312 |
@@
Swin-UMamba | 60M | 68.0G | 0.7760 | 0.8421 |
[…]
@@ 3.4 Comparisons on AbdomenMRI dataset
Table 1 presents the segmentation performance on the AbdomenMRI dataset. Both Swin-UMamba and Swin-UMamba outperform all baseline methods, including CNN-based networks, transformer-based networks, and Mamba-based networks. The superior result demonstrates the great potential of the Mamba-based network in medical image segmentation. Notably, all Mamba-based networks outperform CNN-based and transformer-based baselines by at least 1% on both DSC and NSD. Swin-UMamba exhibits a remarkable 1.34% improvement in DSC over U-Mamba_Enc, which is the previous SOTA model on this dataset. As illustrated in Fig. 3a, Swin-UMamba can recognize the shape and type of target organs, whereas baseline methods fail to accurately identify all target regions.
ImageNet-based pretraining plays a crucial role in our experiments, leading to a significant 7.06% improvement in DSC and a notable 7.74% improvement in NSD for Swin-UMamba. Moreover, leveraging ImageNet-based pretraining facilitates faster and more stable training, requiring merely one-tenth of the training iterations compared to baseline methods. A drastic phenomenon is observed with Swin-UMamba. Without ImageNet-based pretraining, Swin-UMamba fails to converge properly on this dataset with default settings. To address this issue, we disable the deep supervision of Swin-UMamba and extend its training epochs to 200. Despite that, Swin-UMamba outperforms all baseline methods when utilizing the ImageNet pretrained weights. This improvement is particularly noteworthy considering that Swin-UMamba has less than half of the network parameters and FLOPs compared to the previous SOTA model U-Mamba.
We also observed a disparity in parameter numbers and FLOPs between Swin-UMamba and Swin-UMamba. This discrepancy is primarily attributed to the CNN-based decoder, as Swin-UMamba and Swin-UMamba share almost identical structures in the encoder part. We opted to retain the CNN-based decoder to evaluate the impact of pretraining for different models, and it did take better results in this dataset.
@@
Swin-UMamba | 60M | 163.6G | 0.5483 | 0.5632 |
Swin-UMamba | 28M | 45.3G | 0.6402 | 0.6547 |
@@
[…]
Swin-UMamba | 28M | 45.3G | 0.6783 | 0.6933 |
@@ 3.5 Comparisons on Endoscopy dataset
Table 2 presents the segmentation performance of each model on the Endoscopy dataset. Swin-UMamba outperforms U-Mamba_Bot over 2.43% on DSC and 2.41% on NSD. The visualized result of Swin-UMamba on Endoscopy is shown in Fig. 3b. Notably, we observed an impressive performance gain of 12.84% on DSC and 12.90% on NSD with the pretrained model for Swin-UMamba. One possible explanation is that the Endoscopy dataset is smaller than the AbdomenMRI dataset, and models are prone to overfitting to the training data. Leveraging the power of a pretrained model is an effective strategy for mitigating overfitting in such small datasets. In addition, we found that Swin-UMamba performs better than Swin-UMamba on this dataset, possibly benefiting from its fewer parameters to avoid overfitting.
@@
Swin-UMamba | 60M | 174.4G | 0.4561 |
Swin-UMamba | 27M | 48.2G | 0.5186 |
@@
[…]
Swin-UMamba | 27M | 48.2G | 0.5982 |
@@ 3.6 Comparisons on Microscopy dataset
Table 3 presents the segmentation performance on the Microscopy dataset. Swin-UMamba and Swin-UMamba continue to outperform all baseline methods by margins ranging from 1.99% to 20.15%. In contrast to previously mentioned datasets, the Microscopy dataset features higher image resolution, fewer samples, and greater visual difference. This imposes greater demands on the model’s capacity for long-range information modeling and data-efficiency. As shown in Fig. 3c, Swin-UMamba can accurately segment target cells while baselines missing some. Moreover, we observe that Swin-UMamba and Swin-UMamba benefit from the ImageNet pretraining by 12.45% and 7.96% respectively. This once again demonstrates the importance of the pretraining especially for small datasets.